x+y=-9
x-y=14
solve bu substitution

Answer by venugopalramana(3286)

x+y=-9 ...X=-9-Y
x-y=14
-9-Y-Y=14
2Y=-9-14=-23
Y=-23/2=-11.5
X=-9+11.5=2.5